OpenStudy (anonymous):

x^2+12y+y^2+3x-12=0 find the center and radius od the circle

OpenStudy (anonymous):

do you know how to complete the square? because you want it to look like \[(x-h)^2+(y+k)^2=r^2\] so you can read off you answer

OpenStudy (anonymous):

what mertsj said

OpenStudy (anonymous):

i completed the square. i got (x+1.5)^2+(y+6)^2=50.25. i don't know if that's right

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

so my rudius is 7.10 and center (-1.5,-6)?

OpenStudy (anonymous):

yes, if that square root is right (it looks right)